OpenStudy (anonymous):

Write the standard equation of the circle with center (-3, 5) and a diameter of 10. A. (x + 3)2 + (y – 5)2 = 25 B. (x + 3)2 + (y – 5)2 = 100 C. (x – 5)2+ (y + 3)2 = 5 D. (x – 3)2 + (y + 5)2 = 85

hartnn (hartnn):

can you find the radius using diameter ?

hartnn (hartnn):

hint : radius = diameter/2

hartnn (hartnn):

and then you could directly use \(\Large (x-h)^2+(y-k)^2 =r^2\) where h,k = centre = -3,5 and r= radius

OpenStudy (anonymous):

so would the answer be A

hartnn (hartnn):

yes! good :)

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome ^_^